# FlexGuard: A Design Space for On-Body Feedback for Injury Prevention in Strength Training

### v1

@@ 5.1.3. Trigger Location
Rather than naming specific muscles, we describe locations in terms of their functional roles to allow for greater generalizability. Four categories emerged, corresponding to Figure 4. The target region was a common focus, such as in the incline dumbbell fly: “If you started overcompensating and using your wrong muscle, it would like identify that muscle and you’d know you don’t want [that]” (S6). Trainers also warned about overuse of compensation regions, where non-target areas begin to take over the movement. The importance of stabilizing regions, particularly the upper back and core, was emphasized across sessions—“We always want to have them tight core. So tighten our core” (S5). Finally, participants pointed to moving regions (e.g., forearms and hands during an incline fly, or the wrists in a bench press) where feedback can guide trajectory directly. Although the muscle groups at these locations are not necessarily being worked, they act as reference points for maintaining correct movement.
@@ 5.2. Feedback Delivery
Feedback Delivery captures how corrective cues are conveyed to the user once a triggering condition is met as defined in Sensing & Triggering. We define this not only as the onset of feedback, such as vibration, resistance, or pressure, but also its offset. Participants emphasized that the withdrawal of feedback was equally meaningful, since the absence of cues could serve as confirmation that posture had been corrected or movement had returned to the desired path. In this way, Feedback Delivery encompasses both the activation and the release of feedback, as well as its modulation in intensity and timing.
@@ 5.2.1. Feedback Type
extrinsicfeedback can be tuned to produce different actuation patterns. Participants described preferences for continuous cues that guide movement throughout a repetition—“Imagine something helping draw a line for your hands to follow” (N2)—as well as discrete cues that act as one-time warnings when posture falters—“You shouldn’t nag them, just nudge once if they’re falling” (E1).
intrinsicfeedback depends on how materials constrain or guide motion. We distinguish between rigid and elastic responses. rigid feedback provides firm constraints, such as straps that enforce strict ROM limits: “I could feel when I pulled wrong because the strap caught me” (S5). In contrast, elastic feedback allows more flexibility, offering resistance that increases with deviation and guiding users back on track without completely halting movement. As E7 explained, “So a strap that is going to help them maintain the stability. When I’m pressing it, I don’t want to go up and down… if the strap helps us to do that, then it’s pretty good.”
@@ 5.2.2. Feedback Intensity
This dimension reflects how forceful the feedback feels, ranging from low to high. low intensity was favored for subtle prompts or early-stage correction. As E6 explained, “It should whisper, not shout.” In contrast, high intensity was valued for safety-critical moments, such as E4’s proposal for strong restricting feedback in the return phase of a lateral raise: “By the time they’re lowering, it better hit hard if they drop wrong.” Participants also emphasized progressive intensity—feedback that begins subtly and escalates if the error persists or worsens. This approach mirrors how a trainer might tolerate small deviations but intervene more forcefully over time as errors become more critical.
@@ 5.2.3. Feedback Latency
This dimension distinguishes between instant and delayed feedback, reflecting how quickly the system should respond once a deviation is detected. Immediacy is critical in situations involving high injury risk, where feedback must intervene before damage occurs. At the same time, participants recognized the value of delayed feedback that tolerates minor deviations and only escalates if errors persist. Such delayed feedback was seen as less disruptive and more supportive, as it felt less intimidating to users. Several also expressed preferences for progressive onset, where feedback begins subtly and intensifies if the deviation continues. Taken together, these insights suggest that trigger policies must balance the safety benefits of instant correction with the patience of delayed feedback that reduces cognitive overload.
@@ 5.2.4. Feedback Location
This dimension describes where on the body feedback is delivered. From the co-design workshops, we derived body zone maps that highlight regions associated with feedback (Figure 8). extrinsic feedback can be applied directly to the same muscle group identified by Sensing & Triggering, or redirected to a different site to signal correction. In contrast, intrinsic feedback is delivered at the site of motion itself, where resistance provides direct guidance. For example, constraining elbow flare or stabilizing the shoulders.
@@ 5.2.5. Feedback Direction
This dimension describes how corrective forces or cues are oriented relative to the body. We identified four strategies. opposing feedback resists unwanted momentum, such as adding friction in a lateral raise to counter swinging. guiding feedback nudges the user toward the correct path, for example assisting with rep completion or steering the arms back into the proper motion plane, as N6 described when correcting form during a dumbbell incline fly. attentional feedback provides localized cues without displacement, such as vibration in a dumbbell front hold to warn when the arms drop below a safe angle as suggested by N1. Finally, compressive feedback applies inward pressure, for instance stabilizing the core to maintain posture, which participants (e.g., E3, E4, E5) emphasized with their trainees, “always engage the core.”
@@ 5.3. Connecting Sensing & Feedback
The two dimensions of Sensing & Triggering and Feedback Delivery do not operate independently. Rather, they form a closed loop in which detected cues directly shape how feedback is delivered. Figure 10 illustrates this integration across five representative exercises surfaced during co-design workshops, with example approaches for addressing issues at specific Trigger Locations to reduce injury risk. Each row depicts an exercise across its phases of motion, highlighting the Trigger Metrics and Trigger Policies that determine when feedback is activated, alongside the corresponding feedback strategies.
This mapping underscores how expert-based policies can guide the monitoring of metrics (e.g., joint angle, motion path) at locations where injury risk is high. It also extends to personalized policies, such as monitoring muscle activity, which can vary across body types and help detect cases where the intended muscle is not engaged. Once a cue is flagged, Feedback Delivery defines the type and form of response. For example, in a lateral raise, exceeding a speed threshold of movement dynamics triggered intrinsic resistance to slow momentum, which could escalate with progressive intensity, from low to high intensity, if the deviation worsens.
@@ 6. Speed Dating
To evaluate the FlexGuard design space before investing in implementation, we conducted a speed dating workshop. Following prior work using this methodology (Davidoff et al., 2007; Zhou et al., 2023; Lee et al., 2024), we presented a set of storyboards to participants with the same background from which the design space was derived. Each storyboard depicted speculative scenarios combining different feedback modalities, body locations, exercise types, sensing conditions, etc. The goal was to validate and refine the design space with the same user group from which FlexGuard was originally derived, while surfacing the strengths and limitations of different design ideas across scenarios.
We developed ten storyboard scenarios (S1-S10) that instantiated different combinations of Sensing & Triggering and Feedback Delivery dimensions (see Fig 11).
@@ 6.0.1. Study Protocol
We recruited seven beginners (B1-B7) (fewer than three months of consistent practice) to participate in three workshops (ages 24–34; 4 male, 3 female: 28M, 24M, 27F; 26M, 34M; 24M, 25F). Each workshop was conducted in our lab space, which was equipped with two pairs of dumbbells (2 and 12 pounds) and an adjustable workout bench. To establish context, participants were first asked to perform two exercises, dumbbell incline fly and lateral raise, by following short tutorial videos111https://www.youtube.com/@nutritioneering, as these movements cover two primary degrees of freedom of the shoulder. Participants were also free to try additional exercises from the tutorial set if desired. Following this exercise phase, participants were presented with a series of storyboards illustrating scenarios from our design space. They were asked to reflect on whether the scenarios aligned with their needs, expectations, and perceived challenges in strength training. Participants could continue to access the equipment and tutorial videos throughout the session. Each workshop lasted approximately one hour, with the first 10 minutes dedicated to exercise practice and the remaining 50 minutes to storyboard discussion. Participants received $20 in compensation for their time.
@@ 6.1. Results
Two researchers analyzed the workshop transcripts. The analysis surfaced three main themes that validate and extend the FlexGuard design space: (1) semantics of feedback modalities, (2) expectations for timing and progressivity of feedback, and (3) the importance of adaptability across users, exercises, and contexts. These themes highlight how participants navigated trade-offs across the storyboards. Other insights are reported together in a separate group below.
@@ 6.1.1. Semantics of Feedback Modalities
Participants emphasized that different feedback modalities carried distinct meanings in addition to their mechanical effects. intrinsic mechanisms, such as shown in S7, S8 and S9, were perceived as intuitive posture correction. As B2 noted when discussing S8, “If the material itself keeps me from going further, I immediately know what I’m doing wrong.”
extrinsiccues were instead interpreted symbolically: vibration often conveyed “something is wrong or needs attention,” prompting the user to stop and reflect. For example, in bench press (S10), if the right hand drifted ahead of the left, vibration on the right arm could signal the user to pause and allow balance to be restored before continuing. Similarly, pressure applied to a muscle group in S3 resembled a trainer’s tap to activate the correct muscle.
Inconsistent use of feedback semantics risked causing confusion, as B7 emphasized: “The feedback has to be consistent. Otherwise I won’t know what to do.” Several participants envisioned hybrid approaches that combined intrinsic mechanisms for correction with extrinsic cues for reminders or motivational prompts, highlighting the importance of designing semantics for on-body feedback.
6.1.2. Feedback Intensity Spectrum
intensityshaped how participants interpreted and responded to the feedback, with preferences spanning an entire spectrum rather than fixed levels. At the low end, participants described subtle prompts as sufficient for warnings or minor corrections where they could adjust independently. For example, in S3, participants emphasized that feedback should be gentle rather than disruptive: “It should be subtle so it doesn’t make me jump mid-movement, and once I’ve corrected it should go away” At the high end of the spectrum, participants valued strong interventions when safety or performance required it. In S5, participants appreciated the idea of feedback that actively helps complete repetitions: “I’d like this feature—it’s similar to how peers spot each other” (B5). Strong feedback was also desired for critical events such as enforcing safe ROM, as one participant noted in S8: “If I’m about to hurt myself, I want it to stop me right away.” Connecting to the concept of intrinsic feedback, B5 reflected on S7, noting how the intensity spectrum can be expressed through material properties: rigid straps provide high-intensity constraints, while elastic straps offer low-intensity guidance (Fig. 12).
@@ 6.1.3. Adaptability Across Contexts
The results on feedback semantics and intensity clearly underscore the importance of adaptability across different contexts of use. For instance, when performing an exercise, participants expressed a desire for feedback that remains subtle during movement but becomes progressively stronger as the motion deviates, favoring elastic feedback for flexibility (S7). At extrema positions, however, they preferred stronger constraints with rigid feedback to enforce safe ROM (S8). This illustrates the need for wearables that adapt not only across exercises, but also across phases of movement.
Policies such as expert-based can guide these transitions. As in the example above, knowing the thresholds for limiting ROM at both ends allows the system to escalate from elastic to rigid feedback once a defined boundary is reached. Systems must also adapt to users through personalized policies. For example, when fatigue is detected, the system can provide supportive assistance (S5). In S3, if the exercise changes and the target region shifts, the system should adapt reminders to activate the appropriate muscle group. Adaptability also extends to user expertise: as one participant noted, “Once I’ve learned and memorized the correct movement, a reminder might be less necessary.”
@@ 6.1.4. Other Findings
Some scenarios, such as S2, were seen as potentially useful in specific contexts. For example, when training with a mild injury, stabilization could help bear the load from the affected muscles.
At the same time, participants reflected on how wearables might influence learning. Several worried that excessive support could reduce opportunities to internalize correct form. For instance, participants raised concerns about S2: on one hand, users should learn through practice, but on the other, they appreciated how the wearable could provide stabilization to ensure the correct muscles were engaged. In contrast, S9 was regarded as more supportive of learning by providing feedback on tempo. As B1 noted, “I could know if I go too fast or not, because I think that’s important to know.”
Participants also expressed concerns about potential risks with locking mechanisms such as in S6. If weights were falling and users were unable to reach a safe position while locked, this could be dangerous. As B5 explained, “I will hold myself accountable and not rush through the steps, so I might not use it.” To address this, B3 suggested an “emergency button” that would release the lock when things go wrong.
@@ 6.1.5. Summary
Together, these results validate the design space dimensions presented in FlexGuard while also surfacing opportunities for refinement. Because the speed dating method relied on storyboards, participants’ reflections were necessarily speculative rather than grounded in embodied practice despite trying a few exercises in the lab. Nevertheless, adaptability emerged as a recurring priority, particularly for feedback that adjusts across phases of movement. The workshops also clarified the scope of what should be included within the design space. For example, we were initially uncertain whether feedback was necessary for inter-limb coordination (S10), since users can naturally perceive weight imbalance during an exercise. However, across all sessions, participants (B1, B2, B5, B7) consistently emphasized the value of on-body cues to confirm user’s interpretation of the situation—B5: “Although I can feel something is off, the vibration tells me exactly what’s happening.”

### v2

@@ 5.1.3. Trigger Location
Rather than naming specific muscles, we describe locations in terms of their functional roles to allow for greater generalizability. Four categories emerged with a mean of 7 occurences by participants on each, corresponding to Figure 5. The target region was a common focus, such as in the incline dumbbell fly: “If you started overcompensating and using your wrong muscle, it would like identify that muscle and you’d know you don’t want [that]” (S6). Trainers also warned about overuse of compensation regions, where non-target areas begin to take over the movement, as noted by E6 ”If you feel an ache in another area, then you know that you’re not doing it correctly”. The importance of stabilizing regions, particularly the upper back and core, was emphasized across sessions—“If it has a baseline measurement for not engaging your core and then engaging your core, you’ll able to find your center.” (S5). Finally, participants pointed to moving regions (e.g., forearms and hands during an incline fly, or the wrists in a bench press) where feedback can guide trajectory directly. Although the muscle groups at these locations are not necessarily being worked, they act as reference points for maintaining correct movement.
@@ 5.2. Feedback Delivery
Feedback Delivery (Figure 8) captures how corrective cues are conveyed to the user once a triggering condition is met as defined in Sensing & Triggering. We define this not only as the onset of feedback, such as vibration, resistance, or pressure, but also its offset. Participants emphasized that the withdrawal of feedback was equally meaningful, since the absence of cues could serve as confirmation that posture had been corrected or movement had returned to the desired path. In this way, Feedback Delivery encompasses both the activation and the release of feedback, as well as its modulation in intensity and timing.
@@ 5.2.1. Feedback Type
extrinsicfeedback can be tuned to produce different actuation patterns. Participants described preferences for continuous cues that guide movement throughout a repetition—“Imagine something helping draw a line for your hands to follow” (N2) which is a preference raised in 3 out of 9 sessions—as well as discrete cues that act as one-time warnings when posture falters—“You shouldn’t nag them, just nudge once if they’re falling” (E1), emerged in 2 sessions (S1, S3).
intrinsicfeedback depends on how materials constrain or guide motion. We distinguish between rigid and elastic responses. rigid feedback provides firm constraints, such as straps that enforce strict ROM limits: “I could feel when I pulled wrong because the strap caught me” (S5). In contrast, elastic feedback allows more flexibility, offering resistance that increases with deviation and guiding users back on track without completely halting movement. As E7 explained, “So a strap that is going to help them maintain the stability. When I’m pressing it, I don’t want to go up and down… if the strap helps us to do that, then it’s pretty good.” Another participant added ”Elastic is more flexible…it wouldn’t stop me from doing the wrong movement. It would just notify me and I could readjust to my own.” Comments favoring rigid constraints appeared at 5 sessions, while preferences for elastic guidance were raised in 3 sessions.
@@ 5.2.2. Feedback Intensity
This dimension reflects how forceful the feedback feels, ranging from low to high. low intensity was favored for subtle prompts or early-stage correction (mentioned in 3 sessions). As E3 explains to N3, “When someone exercises, they will not have concentration. When it comes to shaking (haptic feedback), it’s not practical.” In contrast, high intensity was valued for safety-critical moments, such as E6’s proposal for strong restricting feedback in the return phase of a lateral raise: “If you started overcompensating and using wrong muscles, it would like identify that muscle and shift your arm back to make sure you’re not using the wrong muscle. An ache in another area would let you know that you’re not doing it correctly.” Participants also emphasized progressive intensity from at least 3 sessions (feedback that begins subtly and escalates if the error persists or worsens) as in S6 where participant highlighted ”Like slowly over time you can do less and less with the pressure.”. This approach mirrors how a trainer might tolerate small deviations but intervene more forcefully over time as errors become more critical.
@@ 5.2.3. Feedback Latency
This dimension distinguishes between instant and delayed feedback, reflecting how quickly the system should respond once a deviation is detected. Immediacy is critical in situations involving high safety risk, where feedback must intervene before damage occurs, ”If the strap helps make sure I don’t go too much in front or too much back… that prevents shoulder injury.” Four sessions emphasized the need for immediate feedback especially for shoulder and elbow. At the same time, participants recognized the value of delayed feedback that tolerates minor deviations and only escalates if errors persist. In S8, trainer states the need for guidance based on tempo not not urgent correction, ”Maybe like click click click… you’d know you’re in the correct position… help with timing and slowing movement down.” Such delayed feedback was seen as less disruptive and more supportive, as it felt less intimidating to users. Several also expressed preferences for progressive onset, where feedback begins subtly and intensifies if the deviation continues. A more tolerable or delayed feedback were discussed in three sessions. Taken together, these insights suggest that Trigger Policies must balance the safety benefits of instant correction with the patience of delayed feedback that reduces cognitive overload.
@@ 5.2.4. Feedback Location
This dimension describes where on the body feedback is delivered. From the co-design workshops, we derived body zone maps that highlight regions associated with feedback (Figure 9). extrinsic feedback can be applied directly to the same muscle group identified by Sensing & Triggering, or redirected to a different site to signal correction. In S5, participant remarks ”Bench press on wrist is the only one I think uses resistance.” A participant also annotated ”A strap in the elbow bend would help” from the co-design workshop in the dumbbell incline fly worksheet as shown in Figure 3. In contrast, intrinsic feedback is delivered at the site of motion itself, where resistance provides direct guidance. For example, constraining elbow flare or stabilizing the shoulders. Several participants highlighted how resistance at the limb could guide technique, such as S4’s observation that “If you tighten it from this point, there’s a chance that I won’t be able to drop my arm down,” underscoring how intrinsic feedback constrains movement at the source.
@@ 5.2.5. Feedback Direction
This dimension describes how corrective forces or cues are oriented relative to the body. We identified four strategies. opposing feedback was discussed in 6 of 9 sessions, to resist unwanted momentum, such as adding friction in a lateral raise to counter swinging as N3 marks ”Right now I feel the restriction in the different direction. When I try to lean my arm toward my body, I feel the restriction this way.”. 4 sessions mentioned guiding feedback nudges the user toward the correct path, for example assisting with rep completion or steering the arms back into the proper motion plane, as N6 described when correcting form during a dumbbell incline fly, ”you know, to kind of like shift your arm back to make sure you’re not using the wrong muscle.”. attentional feedback provides localized cues without displacement, such as vibration in a dumbbell front hold to warn when the arms drop below a safe angle as suggested by N1. Finally, compressive feedback applies inward pressure, for instance stabilizing the core to maintain posture, which participants in three sessions (e.g., E3, E4, E5) emphasized with their trainees, “always engage the core.”
@@ 5.3. Connecting Sensing & Feedback
The two dimensions of Sensing & Triggering and Feedback Delivery do not operate independently. Rather, they form a closed loop in which detected cues directly shape how feedback is delivered. Figure 10 illustrates this integration across five representative exercises surfaced during co-design workshops, with example approaches for addressing issues at specific Trigger Locations to scaffold safety during movement. Each row depicts an exercise across its phases of motion, highlighting the Trigger Metrics and Trigger Policies that determine when feedback is activated, alongside the corresponding feedback strategies.
This mapping underscores how expert-based policies can guide the monitoring of metrics (e.g., joint angle, motion path) at locations where safety risk is elevated. It also extends to personalized policies, such as monitoring muscle activity, which can vary across body types and help detect cases where the intended muscle is not engaged. Once a cue is flagged, Feedback Delivery defines the type and form of response. For example, in a lateral raise, exceeding a speed threshold of movement dynamics triggered intrinsic resistance to slow momentum, which could escalate with progressive intensity, from low to high intensity, if the deviation worsens.
@@ 6. SPEED DATING
To evaluate the FlexGuard design space before investing in implementation, we conducted a speed dating workshop. Following prior work using this methodology (Davidoff et al., 2007; Zhou et al., 2023; Lee et al., 2024), we presented a set of storyboards to participants with the same background from which the design space was derived. Each storyboard depicted speculative scenarios combining different feedback modalities, body locations, exercise types, and sensing conditions. The goal was to validate and refine the design space with users, while surfacing the strengths and limitations of different design ideas across scenarios.
We developed ten storyboard scenarios (S1–S10) that instantiated different combinations of Sensing & Triggering and Feedback Delivery dimensions (see Fig. 11).
@@ 6.0.1. Study Protocol
We recruited seven beginners (B1–B7) with fewer than three months of consistent strength training experience to participate in three workshops (ages 24–34; 4 male, 3 female). Each workshop was conducted in our lab space, which was equipped with two pairs of dumbbells (2 and 12 pounds) and an adjustable workout bench. To establish context, participants were first asked to perform two exercises, dumbbell incline fly and lateral raise, by following short tutorial videos111https://www.youtube.com/@nutritioneering, as these movements cover two primary degrees of freedom of the shoulder. Participants were also free to try additional exercises from the tutorial set if desired.
Following this exercise phase, participants were presented with a series of storyboards illustrating scenarios from the FlexGuard design space. They were asked to reflect on whether the scenarios aligned with their needs, expectations, and perceived challenges in strength training. Participants could continue to access the equipment and tutorial videos throughout the session. Each workshop lasted approximately one hour, with the first 10 minutes dedicated to exercise practice and the remaining 50 minutes to storyboard discussion. Participants received $20 in compensation for their time.
@@ 6.1. Results
Two researchers analyzed the workshop transcripts. This analysis clarified and validated our assumptions about FlexGuard, allowing us to refine the proposed design space. In addition, we identified recurring themes that reflect underlying user needs and design considerations emerging from the speculative storyboard scenarios. These insights inform the design of future on-body feedback systems aimed at scaffolding safety during strength training, as participants reasoned about both becoming aware of safety-relevant movement and feeling physically supported during execution.
@@ 6.1.1. Semantics of Feedback Modalities
Participants suggested that different feedback modalities could carry distinct meanings beyond their mechanical effects. intrinsic mechanisms, such as those shown in S7, S8, and S9, were commonly perceived as intuitive cues for posture correction, as reflected in 5 out of 7 participants’ interpretations. As B2 noted when discussing S8, “If the material itself keeps me from going further, I immediately know what I’m doing wrong.”
In contrast, extrinsic cues were interpreted more symbolically. Vibration often conveyed that “something is wrong or needs attention,” prompting users to pause and reflect. For example, in bench press (S10), if the right hand drifted ahead of the left, vibration on the right arm could signal the user to pause and allow balance to be restored before continuing. Similarly, pressure applied to a muscle group in S3 was described as resembling a trainer’s tap to draw attention to the correct muscle.
Participants also speculated that inconsistent feedback semantics could cause confusion, as B7 remarked: “The feedback has to be consistent. Otherwise I won’t know what to do.” Several participants envisioned hybrid approaches that combine intrinsic mechanisms for corrective guidance with extrinsic cues for reminders or motivational prompts, highlighting the importance of consistent and interpretable semantics for scaffolding safety.
6.1.2. Movement Continuity
Participants frequently reasoned about on-body feedback in terms of whether it allowed them to move smoothly through an exercise or interrupted movement at critical moments. Rather than focusing on specific modalities or implementations, the feedback preferences were around maintaining movement continuity during safe execution and accepting interruption when safety risks increased (Figure 12).
During continuous movement, participants preferred feedback that minimally interfered with execution, allowing them to maintain rhythm and comfort while remaining aware of deviations. In these phases, lower levels of interference were favored, as participants felt able to adjust independently without disruption. For example, in S3, participants emphasized that feedback should remain subtle during execution: “It should be subtle so it doesn’t make me jump mid-movement, and once I’ve corrected it should go away.” Participants also suggested that feedback could relax or reset after the execution phase, reflecting a desire to avoid persistent interruption when movement was perceived as safe.
In contrast, when participants perceived heightened safety risk, such as fatigue-induced shaking or approaching range-of-motion boundaries, they valued feedback that interrupted movement continuity. In S5, participants appreciated feedback that provided physical assistance to help complete remaining repetitions while making limits explicit: “I’d like this feature—it’s similar to how peers spot each other” (B5). Similarly, in S8, participants described the usefulness of feedback that made further movement difficult at boundaries: “If I’m about to hurt myself, I want it to stop me right away.”
@@ 6.1.3. Adaptability Across Contexts
Participants’ reflections on feedback semantics and intensity suggested that adaptability is important across different contexts of use. For instance, when performing an exercise, participants preferred feedback that remains subtle during movement but becomes progressively stronger as deviations increase, favoring elastic feedback for flexibility (S7). At extrema positions, however, they preferred stronger constraints with rigid feedback to clearly signal safe ROM boundaries (S8). This indicates that participants imagined wearables that adapt not only across exercises, but also across different phases of movement.
Policies such as expert-based were mentioned as one way to guide such transitions. Participants imagined systems that could use thresholds for limiting ROM at both ends to escalate from elastic to rigid feedback once a boundary is reached (mentioned by B1, B2, B5, and B6). Participants also speculated about personalized policies, such as adapting feedback based on fatigue (S5) or changing reminders as the exercise and target region shift. Adaptability was also expected to vary with user expertise. For instance, one participant noted, “Once I’ve learned and memorized the correct movement, a reminder might be less necessary.”
@@ 6.1.4. Other Findings
Some scenarios, such as S2, were seen as useful in specific contexts. For example, when training with a mild injury, stabilization could help redistribute load away from affected muscles.
At the same time, participants reflected on how wearables might influence learning. Several expressed concern that excessive support could reduce opportunities to internalize correct form. For instance, in S2, participants noted a tension between learning through practice and appreciating stabilization that helps ensure correct muscle engagement. In contrast, S9 was regarded as more supportive of learning by providing feedback on tempo. As B1 noted, “I could know if I go too fast or not, because I think that’s important to know.”
Participants also raised concerns about potential risks associated with locking mechanisms such as in S6. If weights were falling and users were unable to reach a safe position while locked, this could be dangerous. As B5 explained, “I will hold myself accountable and not rush through the steps, so I might not use it.” To address this concern, B3 suggested including an “emergency button” that would release the lock when needed.
@@ 6.1.5. Summary
Together, these results validate the dimensions of the FlexGuard design space while also surfacing opportunities for refinement. Because the speed dating method relied on storyboards, participants’ reflections were necessarily speculative rather than grounded in embodied practice despite trying a few exercises in the lab. Nevertheless, adaptability emerged as a recurring priority, particularly for feedback that adjusts across phases of movement. The workshops also clarified the scope of what should be included within the design space. For example, we were initially uncertain whether feedback was necessary for inter-limb coordination (S10), since users can naturally perceive weight imbalance during an exercise. However, across all sessions, participants (B1, B2, B5, B7) consistently emphasized the value of on-body cues to confirm user’s interpretation of the situation—B5: “Although I can feel something is off, the vibration tells me exactly what’s happening.”